# The Best of Both Worlds: Integrating Language Models and Diffusion Models for Video Generation

### v1

@@ Abstract
Recent advancements in text-to-video (T2V) generation have been driven by two competing paradigms: autoregressive language models and diffusion models. However, each paradigm has intrinsic limitations: language models struggle with visual quality and error accumulation, while diffusion models lack semantic understanding and causal modeling. In this work, we propose LanDiff, a hybrid framework that synergizes the strengths of both paradigms through coarse-to-fine generation. Our architecture introduces three key innovations: (1) a semantic tokenizer that compresses 3D visual features into compact 1D discrete representations through efficient semantic compression, achieving a $\sim$14,000$\times$ compression ratio; (2) a language model that generates semantic tokens with high-level semantic relationships; (3) a streaming diffusion model that refines coarse semantics into high-fidelity videos. Experiments show that LanDiff, a 5B model, achieves a score of 85.43 on the VBench T2V benchmark, surpassing the state-of-the-art open-source models Hunyuan Video (13B) and other commercial models such as Sora, Keling, and Hailuo. Furthermore, our model also achieves state-of-the-art performance in long video generation, surpassing other open-source models in this field. Our demo can be viewed at https://landiff.github.io/.
@@ 1 Introduction
Thanks to these designs, our LanDiff has made significant progress in spatial relationship compliance, action coherence, visual quality, etc. Specifically, 1) for short video generation, our 5B model achieved a score of 85.43 on the VBench T2V benchmark, surpassing the state-of-the-art open-source models Hunyuan Video (13B) and other commercial models such as Sora, Keling, and Hailuo; 2) for long video generation, after testing by the VBench T2V benchmark, our model also achieved state-of-the-art performance, surpassing other open-source models in this field. Our video examples can be viewed at https://landiff.github.io/.
@@ 2.1 Video Semantic Tokenizer
Based on these thoughts, we choose the pretrained SSL features as the video representations. For better usage of different SSL models, we choose Theia model (Shang et al., 2024), which is a unified visual feature extractor distilled from multiple visual task models, including CLIP (Radford et al., 2021), (Kirillov et al., 2023), DINOv2 (Oquab et al., 2024), ViT (Dosovitskiy et al., 2021), and Depth-anything (Yang et al., 2024a).
@@ 2.3 Diffusion Model for Perceptual Feature Generation
3 Experiments and Results
3.1 Experimental Settings
@@
3.2 Experimental Results
Text to Video Generation. As shown in Table 3, we conducted a quantitative comparison between LanDiff and other state-of-the-art open-source models on the VBench benchmark. In the table, we selected 9 dimensions related to temporal modeling and semantics for presentation, and we show the scores of all 16 dimensions in the appendix. Our 5B LanDiff model outperforms the state-of-the-art open-source 13B Hunyuan Video model and achieves a score of 85.42. Our model achieves the highest semantic score and the second-highest quality score, indicating that our model can follow the text well and generate high-quality videos. Compared with the same-sized CogVideoX-5B, our model achieves better results in almost all dimensions. To eliminate the interference of the training data volume, we train a 7B Dit-based diffusion model with the same training data recipe as LanDiff. We mark it as Dit in the table. This indicates that with the same data volume, by reasonably dividing the task load, we can combine the advantages of autoregressive models and diffusion models to achieve better performance than diffusion models alone. Figure 6 shows the qualitative comparison between the videos generated by LanDiff and CogVideoX-5B. In the first example, it can be seen that in the video generated by CogVideoX-5B, one fish disappears after the fish meet, while in the video generated by our model, the fish can still remain intact after meeting. This demonstrates that our model can understand the concept of fish as an entity and maintain consistency over time. Additionally, our model more faithfully adheres to the background elements described in the prompt “with rocks and shadows in the background”, rendering these environmental details with greater accuracy and consistency. In the second example, our model accurately captures the temporal dynamics described in the prompt, successfully rendering a melting ice sculpture of a dog. LanDiff properly depicts the progressive melting process, with the ice dog gradually losing its form and eventually transforming into a puddle of water. In contrast, CogVideoX-5B generates a static representation of the ice dog sculpture that remains largely unchanged throughout the sequence, failing to capture the crucial temporal narrative of melting described in the prompt. This highlights our model’s superior capability in understanding and visualizing complex temporal transformations and physical processes.
Long Video Generation. We compare our model with other open-source text-to-long video generation models: 1) FreeNoise (Qiu et al., 2024). 2) StreamingT2V (Henschel et al., 2024). 3) OpenSora-V1.2 (Zheng et al., 2024). 4) ARLON (Li et al., 2024). As shown in Table 4, our LanDiff achieves state-of-the-art performance in long video generation tasks with the highest Total Score of 68.34, outperforming other open-source models across almost all dimensions. While StreamingT2V exhibits the highest Dynamic Degree (85.64), this comes at a clear cost to consistency metrics, as evidenced by its significantly lower Subject Consistency (87.31) and Background Consistency (94.64). An optimal video generation model should maximize dynamic content while maintaining temporal coherence. Compared with models that excel in consistency metrics but show limited dynamism (such as ARLON with 50.42 Dynamic Degree), our approach demonstrates superior dynamism (72.86) while preserving strong consistency scores and achieving the best results in Aesthetic Quality (60.96) and Overall Consistency (27.29). As illustrated in Figure 5, LanDiff effectively generates complex motion dynamics when prompted with “a car accelerating to gain speed,” successfully depicting the progressive increase in vehicle velocity while preserving visual consistency across frames. In contrast, FreeNoise, StreamingT2V, and OpenSora-V1.2 all struggle to properly render the acceleration motion, exhibiting either static vehicles, inconsistent car appearances, or unrealistic motion patterns that fail to convey the sense of increasing speed.
@@ 3.2 Experimental Results
Video Tokenization. We present visualization results of our model’s video reconstruction in Figure 7. We extract semantic tokens from the reference video and subsequently convert these tokens back into video using the video detokenizer. The results demonstrate that through our careful design, our video tokenizer can accurately reconstruct videos with accurate semantics and actions using only approximately 1/50th of the sequence length required by Magvit2 (Yu et al., 2024a). While some minor discrepancies in clothing details exist between the reconstructed and reference videos, these differences remain within acceptable limits for practical video generation applications. Figure 8 shows the reconstruction loss trajectories for IFrame and PFrame components during training. Despite allocating significantly different token quantities (330 vs. 74), both frame types converge to comparable reconstruction quality. This validates our video frame grouping strategy that prioritizes key frames while minimizing tokens for intermediate frames. The results confirm our approach successfully balances reconstruction fidelity with computational efficiency.
@@ 4.1 Video Tokenization
Video tokenization plays a crucial role in video understanding and generation tasks. Since a video can be represented as a sequence of continuous frames, some works directly use image tokenizers to process videos frame by frame. For example, Wang et al. (2024a) directly use SBER-MoVQGAN as the video tokenizer. However, this method ignores the temporal redundancy in videos, resulting in a low compression rate. To reduce temporal redundancy, some works (Yan et al., 2021; Yu et al., 2024a) try to extend image tokenizers based on 2D convolution to 3D convolution, which can process both temporal and spatial information simultaneously. These methods encode videos in the original RGB space and perform video reconstruction tasks, which are more about perceptual compression. In addition, some works (Ge et al., 2023; Jin et al., 2023) try to train video tokenizers on features extracted from pre-trained visual encoders. These tokenizers can achieve good performance in understanding and generation tasks while maintaining a high compression rate. The video tokenizer we propose belongs to this category. Unlike previous feature-based tokenizers that achieved limited compression rates and primarily focused on image processing, our video tokenizer delivers significantly higher compression while handling both images and videos in a unified framework.
4.2 LLM based Video Generation
@@
LLM-based video generation methods are usually based on the Transformer (Vaswani et al., 2017) structure, learning the mapping from text to video through next token prediction. TATS (Ge et al., 2022) uses VQ-GAN as the tokenizer and predicts video tokens using a GPT-like model structure. VideoGPT (Yan et al., 2021) uses 3D convolution to extract features and quantize them, and predicts the quantized video discrete tokens using a GPT-like model. Recently, VideoPoet (Yu et al., 2024a) uses MagViT2 (Yu et al., 2024a) as the tokenizer and unifies multiple modalities as input to a large language model (LLM) to conditionally generate video tokens. In addition, Emu3 (Wang et al., 2024a) uses SBER-MoVQGAN as the tokenizer to perform video understanding and generation by predicting the next token. These works all use LLMs to directly generate “perceptual features” that contain rich visual details with high bit rates. Recently, ARLON (Li et al., 2024) attempts to discretize VAE features into a small number of tokens to reduce the bit rate required for LLMs’ prediction. In this way, the tokens retain low-frequency visual information such as blurry contours rather than high-level semantic information. In contrast, our method employs tokens containing high-level semantic information as prediction targets for LLMs, which enables us to fully leverage LLMs’ advantages in causal modeling to precisely generate high-quality videos.
4.3 Diffusion based Video Generation
@@
Diffusion-based methods have achieved great success in image generation, and recently many people (Chen et al., 2024; Ho et al., 2022a, b; Singer et al., 2023; Zhou et al., 2023) have tried to apply them to video generation tasks. VDM (Ho et al., 2022b) extends the 3D U-Net structure for video generation. Wang et al. (2024b) propose to generate high-quality and aesthetically pleasing videos in a cascaded manner. Benefiting from the success of the text-to-image (T2I) field, some works such as Animatediff (Guo et al., 2024), SVD (Blattmann et al., 2023), and PixelDance (Zeng et al., 2023) try to use pre-trained T2I models as initialization, and then add modules for temporal modeling to capture motion information for video generation. Ma et al. (2024) explore the generation capabilities of multiple different structures of latent diffusion transformer. After the release of SORA, a series of video generation methods based on the DiT (Peebles & Xie, 2023) model have been proposed, including OpenSora (Zheng et al., 2024), OpenSoraPlan (Lin et al., 2024a), Cogvideox (Yang et al., 2024b), Hunyuan Video (Kong et al., 2024), Mira (Ju et al., 2024) and STIV (Lin et al., 2024b) etc. These methods can only generate short videos of a few seconds. Recently, StreamingT2V (Henschel et al., 2024) generates long videos by block-wise generation on a pre-trained short video generation model, and then uniformly performs mixed augmentation. In addition, some works improve the consistency of long video generation by leveraging noise rescheduling techniques (Qiu et al., 2024; Lu et al., 2024). Our method employs diffusion models as renderers for semantic features, enabling us to leverage their superior visual generation quality while circumventing their limitations in causal modeling.
@@ Appendix A Implementation Details
Total 170M
A.1 Details of Video Tokenizer
@@
Total 5B
@@ A.3 Details of Diffusion Model
Figure 9: Radar chart visualization of performance comparison across different dimensions on VBench. The plot compares LanDiff against five competitive baselines: Sora, Hailuo, HunyuanVideo, Kling, and CogVideoX-5B.
A.4 More Analysis on VBench Benchmark
@@
As visualized in Figure 9, we compare LanDiff with five representative models across different evaluation dimensions. The radar chart reveals that LanDiff (shown in red) demonstrates well-balanced performance across all metrics, with notably strong results in quality score and semantic accuracy. While Sora (shown in blue) achieves competitive scores in imaging quality and scene, and HunyuanVideo excels in certain visual aspects, LanDiff maintains consistently superior performance across the entire spectrum of metrics. The comprehensive comparison with these strong baselines, including both commercial (Sora, Hailuo, Kling) and open-source models (HunyuanVideo, CogVideoX-5B), further validates the effectiveness of our hybrid LLM+Diffusion approach.
A.5 More Examples

### v3

@@ Abstract
Recent advancements in text-to-video (T2V) generation have been driven by two competing paradigms: autoregressive language models and diffusion models. However, each paradigm has intrinsic limitations: language models struggle with visual quality and error accumulation, while diffusion models lack semantic understanding and causal modeling. In this work, we propose LanDiff, a hybrid framework that synergizes the strengths of both paradigms through coarse-to-fine generation. Our architecture introduces three key innovations: (1) a semantic tokenizer that compresses 3D visual features into compact 1D discrete representations through efficient semantic compression, achieving a $\sim$14,000$\times$ compression ratio; (2) a language model that generates semantic tokens with high-level semantic relationships; (3) a streaming diffusion model that refines coarse semantics into high-fidelity videos. Experiments show that LanDiff, a 5B model, achieves a score of 85.43 on the VBench T2V benchmark, surpassing the state-of-the-art open-source models Hunyuan Video (13B) and other commercial models such as Sora, Kling, and Hailuo. Furthermore, our model also achieves state-of-the-art performance in long video generation, surpassing other open-source models in this field. Our demo can be viewed at https://landiff.github.io/.
@@ 1 Introduction
Thanks to these designs, our LanDiff has made significant progress in spatial relationship compliance, action coherence, visual quality, etc. Specifically, 1) for short video generation, our 5B model achieved a score of 85.43 on the VBench T2V benchmark, surpassing the state-of-the-art open-source models Hunyuan Video (13B) and other commercial models such as Sora, Kling, and Hailuo; 2) for long video generation, after testing by the VBench T2V benchmark, our model also achieved state-of-the-art performance, surpassing other open-source models in this field. Our video examples can be viewed at https://landiff.github.io/.
@@ 2.1 Video Semantic Tokenizer
Based on these thoughts, we choose the pretrained SSL features as the video representations. For better usage of different SSL models, we choose Theia model (Shang et al., 2024), which is a unified visual feature extractor distilled from multiple visual task models, including CLIP (Radford et al., 2021), SAM (Kirillov et al., 2023), DINOv2 (Oquab et al., 2024), ViT (Dosovitskiy et al., 2021), and Depth-Anything (Yang et al., 2024a).
@@ 2.3 Diffusion Model for Perceptual Feature Generation
3 Experiments and Results
3.1 Experimental Settings
3.2 Experimental Results
@@
Text to Video Generation. As shown in Table 3, we conducted a quantitative comparison between LanDiff and other state-of-the-art open-source models on the VBench benchmark. Our model achieves the highest semantic score and quality score, indicating that our model can follow the text well and generate high-quality videos. Compared with the same-sized CogVideoX-5B, our model achieves better results in almost all dimensions. To eliminate the interference of the training data volume, we train a 7B DiT-based diffusion model with the same training data recipe as LanDiff. We mark it as DiT in the table. This indicates that with the same data volume, by reasonably dividing the task load, we can combine the advantages of autoregressive models and diffusion models to achieve better performance than diffusion models alone. Figure 5 shows the qualitative comparison between the videos generated by LanDiff and CogVideoX-5B. In the first example, it can be seen that in the video generated by CogVideoX-5B, one fish disappears after the fish meet, while in the video generated by our model, the fish can still remain intact after meeting. This demonstrates that our model can understand the concept of fish as an entity and maintain consistency over time. Additionally, our model more faithfully adheres to the background elements described in the prompt “with rocks and shadows in the background”, rendering these environmental details with greater accuracy and consistency. In the second example, our model accurately captures the temporal dynamics described in the prompt, successfully rendering a melting ice sculpture of a dog. LanDiff properly depicts the progressive melting process, with the ice dog gradually losing its form and eventually transforming into a puddle of water. In contrast, CogVideoX-5B generates a static representation of the ice dog sculpture that remains largely unchanged throughout the sequence, failing to capture the crucial temporal narrative of melting described in the prompt. This highlights our model’s superior capability in understanding and visualizing complex temporal transformations and physical processes. For more video examples, please refer to our demo website [^2] https://landiff.github.io/.
@@
Long Video Generation. We compare our model with other open-source text-to-long video generation models: 1) FreeNoise (Qiu et al., 2024). 2) StreamingT2V (Henschel et al., 2024). 3) OpenSora-V1.2 (Zheng et al., 2024). 4) ARLON (Li et al., 2024). As shown in Table 4, our LanDiff achieves state-of-the-art performance in long video generation tasks with the highest Total Score of 68.34, outperforming other open-source models across almost all dimensions. While StreamingT2V exhibits the highest Dynamic Degree (85.64), this comes at a clear cost to consistency metrics, as evidenced by its significantly lower Subject Consistency (87.31) and Background Consistency (94.64). An optimal video generation model should maximize dynamic content while maintaining temporal coherence. Compared with models that excel in consistency metrics but show limited dynamism (such as ARLON with 50.42 Dynamic Degree), our approach demonstrates superior dynamism (72.86) while preserving strong consistency scores and achieving the best results in Aesthetic Quality (60.96) and Overall Consistency (27.29). As illustrated in Figure 6, LanDiff effectively generates complex motion dynamics when prompted with “a car accelerating to gain speed,” successfully depicting the progressive increase in vehicle velocity while preserving visual consistency across frames. In contrast, FreeNoise, StreamingT2V, and OpenSora-V1.2 all struggle to properly render the acceleration motion, exhibiting either static vehicles, inconsistent car appearances, or unrealistic motion patterns that fail to convey the sense of increasing speed. Additional video samples are available on our demo site2.
@@
Video Tokenization. We present visualization results of our model’s video reconstruction in Figure 7. We extract semantic tokens from the reference video and subsequently convert these tokens back into video using the video detokenizer. The results demonstrate that through our careful design, our video tokenizer can accurately reconstruct videos with accurate semantics and actions using only approximately 1/50th of the sequence length required by MagViT2 (Yu et al., 2024a). While some minor discrepancies in clothing details exist between the reconstructed and reference videos, these differences remain within acceptable limits for practical video generation applications. Additional video samples are available on our demo site2. Figure 8 shows the reconstruction loss trajectories for IFrame and PFrame components during training. Despite allocating significantly different token quantities (330 vs. 74), both frame types converge to comparable reconstruction quality. This validates our video frame grouping strategy that prioritizes key frames while minimizing tokens for intermediate frames. The results confirm our approach successfully balances reconstruction fidelity with computational efficiency.
@@ 4 Related Work
Video Tokenization. Video tokenization plays a crucial role in video understanding and generation tasks. Since a video can be represented as a sequence of continuous frames, some works directly use image tokenizers to process videos frame by frame. For example, Wang et al. (2024a) directly use SBER-MoVQGAN as the video tokenizer. However, this method ignores the temporal redundancy in videos, resulting in a low compression rate. To reduce temporal redundancy, some works (Yan et al., 2021; Yu et al., 2024a) try to extend image tokenizers based on 2D convolution to 3D convolution, which can process both temporal and spatial information simultaneously. These methods encode videos in the original RGB space and perform video reconstruction tasks, which are more about perceptual compression. In addition, some works (Ge et al., 2023; Jin et al., 2023) try to train video tokenizers on features extracted from pre-trained visual encoders. These tokenizers can achieve good performance in understanding and generation tasks while maintaining a high compression rate. The video tokenizer we propose belongs to this category. Unlike previous feature-based tokenizers that achieved limited compression rates and primarily focused on image processing, our video tokenizer delivers significantly higher compression while handling both images and videos in a unified framework.
@@
LLM based Video Generation. LLM-based video generation methods are usually based on the Transformer (Vaswani et al., 2017) structure, learning the mapping from text to video through next token prediction. TATS (Ge et al., 2022) uses VQ-GAN as the tokenizer and predicts video tokens using a GPT-like model structure. VideoGPT (Yan et al., 2021) uses 3D convolution to extract features and quantize them, and predicts the quantized video discrete tokens using a GPT-like model. Recently, VideoPoet (Yu et al., 2024a) uses MagViT2 (Yu et al., 2024a) as the tokenizer and unifies multiple modalities as input to a large language model (LLM) to conditionally generate video tokens. In addition, Emu3 (Wang et al., 2024a) uses SBER-MoVQGAN as the tokenizer to perform video understanding and generation by predicting the next token. These works all use LLMs to directly generate “perceptual features” that contain rich visual details with high bit rates. Recently, ARLON (Li et al., 2024) attempts to discretize VAE features into a small number of tokens to reduce the bit rate required for LLMs’ prediction. In this way, the tokens retain low-frequency visual information such as blurry contours rather than high-level semantic information. In contrast, our method employs tokens containing high-level semantic information as prediction targets for LLMs, which enables us to fully leverage LLMs’ advantages in causal modeling to precisely generate high-quality videos.
@@
Diffusion based Video Generation. Diffusion-based methods have achieved great success in image generation, and recently many people (Chen et al., 2024; Ho et al., 2022a, b; Singer et al., 2023; Zhou et al., 2023) have tried to apply them to video generation tasks. VDM (Ho et al., 2022b) extends the 3D U-Net structure for video generation. Wang et al. (2024b) propose to generate high-quality and aesthetically pleasing videos in a cascaded manner. Benefiting from the success of the text-to-image (T2I) field, some works such as Animatediff (Guo et al., 2024), SVD (Blattmann et al., 2023), and PixelDance (Zeng et al., 2023) try to use pre-trained T2I models as initialization, and then add modules for temporal modeling to capture motion information for video generation. Ma et al. (2024) explore the generation capabilities of multiple different structures of latent diffusion transformer. After the release of SORA, a series of video generation methods based on the DiT (Peebles & Xie, 2023) model have been proposed, including OpenSora (Zheng et al., 2024), OpenSoraPlan (Lin et al., 2024a), Cogvideox (Yang et al., 2024b), Hunyuan Video (Kong et al., 2024), Mira (Ju et al., 2024) and STIV (Lin et al., 2024b) etc. These methods can only generate short videos of a few seconds. Recently, StreamingT2V (Henschel et al., 2024) generates long videos by block-wise generation on a pre-trained short video generation model, and then uniformly performs mixed augmentation. In addition, some works improve the consistency of long video generation by leveraging noise rescheduling techniques (Qiu et al., 2024; Lu et al., 2024). Our method employs diffusion models as renderers for semantic features, enabling us to leverage their superior visual generation quality while circumventing their limitations in causal modeling.
@@ 5 Conclusion
6 Broader Impact
Text-to-video generation models such as LanDiff offer substantial potential for creative applications across entertainment, education, and content creation domains. Nevertheless, these technologies introduce ethical considerations and potential risks that warrant attention. The capability to generate photorealistic videos from textual descriptions could potentially be exploited to create misleading or deceptive content, including sophisticated deepfakes or videos that misrepresent individuals or events. Such misuse raises significant concerns regarding misinformation propagation, privacy violations, and potential harm to individuals or communities. To mitigate these risks, we recommend several safeguards: (1) incorporating visible watermarks or robust digital signatures in generated content to ensure transparency regarding its synthetic nature; (2) advancing and deploying sophisticated detection systems capable of identifying AI-generated content with high accuracy; (3) establishing comprehensive usage policies that explicitly prohibit harmful applications; and (4) implementing accessible reporting mechanisms for suspected misuse cases. Furthermore, we advocate for continued research into technical safeguards and responsible deployment frameworks specifically designed for generative video models. We emphasize that our research contribution aims to advance the field of multimodal generation for socially beneficial applications while acknowledging the necessity of addressing potential negative impacts through complementary technical innovations and policy measures.
7 Limitation and Future Works
While LanDiff demonstrates significant advancements in text-to-video generation, several limitations remain to be addressed in future work. First, the scale of our language model (2B parameters) is substantially smaller than state-of-the-art text-only LLMs, potentially limiting the semantic understanding and generation capabilities. Future work will explore scaling our language models to larger parameter counts to enhance performance. Second, our choice of CogVideoX-2B as the underlying diffusion model establishes an upper bound on the quality of generated videos. We plan to investigate the development of more sophisticated diffusion backbones specifically optimized for video generation tasks to overcome this constraint. Third, we observe that LanDiff struggles with accurate text rendering within generated videos. This limitation likely stems from insufficient supervision of text features in the current semantic token representation. Future research will focus on developing more comprehensive video semantic tokens with enhanced text-specific supervision. Our current work primarily addresses text-to-video generation, but we envision extending LanDiff’s capabilities to broader application scenarios. These include image-to-video generation, unified models for video understanding and generation, and interactive controllable video synthesis. These extensions would significantly expand the utility of semantic token-based language models in multimodal generation tasks and provide more flexible creative tools for users across various domains.
@@ Appendix A Implementation Details
Total Parameters 170M
@@
Trainable Parameters 3B
Total Parameters 5B
A.1 Details of Video Tokenizer
@@ A.3 Details of Diffusion Model
Figure 9: Radar chart visualization of performance comparison across different dimensions on VBench. The plot compares LanDiff against five competitive baselines: Sora, Hailuo, HunyuanVideo, Kling, and CogVideoX-5B. For better readability, the values in the radar chart have been normalized to a scale ranging from 0.3 to 0.8. The normalization was performed using the min-max scaling formula: $normalized=0.3+0.5\times\frac{value-min\_value}{max\_value-min\_value}$. The original raw performance data can be found in Table 3 and Table 8.
Appendix B More Analysis on VBench Benchmark
@@
As visualized in Figure 9, we compare LanDiff with five representative models across different evaluation dimensions. The radar chart reveals that LanDiff (shown in red) demonstrates well-balanced performance across all metrics, with notably strong results in quality score and semantic accuracy. While Sora (shown in blue) achieves competitive scores in imaging quality and scene, and HunyuanVideo excels in certain visual aspects, LanDiff maintains consistently superior performance across the entire spectrum of metrics. Notably, while there is typically a trade-off between dynamic degree and motion smoothness/subject consistency, LanDiff achieves a high level of dynamism while maintaining strong performance in stability metrics - with subject consistency and motion smoothness scores within 2.3% and 2.5% of the best-performing models respectively. Notably, while text-to-video models typically exhibit a trade-off between dynamic expressiveness and temporal stability, LanDiff successfully balances these competing objectives—achieving high dynamism while maintaining robust stability metrics, with subject consistency and motion smoothness scores deviating by only 2.3% and 2.5% respectively from the state-of-the-art in each category. The comprehensive comparison with these strong baselines, including both commercial (Sora, Hailuo, Kling) and open-source models (HunyuanVideo, CogVideoX-5B), further validates the effectiveness of our hybrid LLM+Diffusion approach.
Appendix C More Examples